OpenStudy (anonymous):

Mason is making bread from a recipe that calls for 6 4/17 cups of flour. If he increases the recipe 2 2/3 times, about how many cups of flour will Mason need?

OpenStudy (anonymous):

LET X= cups of flour. x=(6 4/17)(2 2/3) x=(106/17)(8/3) x=16 and 32/51 cups of flour PLEASE CLICK THE GOOD ANSWER BUTTON TNX!